# Auditing Google’s AI Overviews and Featured Snippets: A Case Study on Baby Care and Pregnancy

### v2

@@ Abstract
Google Search increasingly surfaces AI-generated content through features like AI Overviews (AIO) and Featured Snippets (FS), which users frequently rely on despite having no control over their presentation. Through a systematic algorithm audit of 1,508 real baby care and pregnancy-related queries, we evaluate the quality and consistency of these information displays. Our robust evaluation framework assesses multiple quality dimensions, including answer consistency, relevance, presence of medical safeguards, source categories, and sentiment alignment. Our results reveal concerning gaps in information consistency, with information in AIO and FS displayed on the same search result page being inconsistent with each other in 33% of cases. Despite high relevance scores, both features critically lack medical safeguards (present in just 11% of AIO and 7% of FS responses). While health and wellness websites dominate source categories for both, AIO and FS, FS also often link to commercial sources. These findings have important implications for public health information access and demonstrate the need for stronger quality controls in AI-mediated health information. Our methodology provides a transferable framework for auditing AI systems across high-stakes domains where information quality directly impacts user well-being.
@@
Figure 1: Overview of our audit study methods and results on Google’s AI Overviews (AIO) and Featured Snippets (FS) for pregnancy and baby care information. In the example screenshot, the AIO appears above the FS. The first sentence of the AOI is highlighted by Google to add emphasis. We find that AIO occur more frequently than FS (RQ1) and show considerable inconsistency with FS answers, more profound in highlighted pairs (RQ2). While AIO and FS responses are generally relevant (RQ3a), they provide safeguard cues infrequently (RQ3b), and FS sources concern significantly higher proportions from commercial categories than AIO or ten blue links results (RQ3c). RQ4: We do not find evidence of “confirmation bias” in AIO answers, where a user’s sentiment (positive/neutral/negative) is reflected in the query.
Introduction
@@
Search engines like Google serve as critical information gatekeepers for high-stakes topics where people need to rapidly acquire extensive knowledge to make important decisions. This includes domains like healthcare, finance, and politics, where search results significantly impact individual and societal well-being (Fox 2011; Fernández-Pichel, Pichel, and Losada 2024). Among these areas, baby care and pregnancy offer a compelling example – new and expectant parents make frequent, time-sensitive decisions while learning from many new concepts, often relying on search engines during a critical period lasting months or years. Specifically, 85% of pregnant women surveyed by (Giacometti et al. 2024) turned to Google for pregnancy-related information in 2024. While over 80% stated that the information they found online had a medium to very high influence on their decision-making relating to pregnancy, only 42% reported validating such information with their doctors (Giacometti et al. 2024). Thus, web search, particularly Google, has become a primary source for crucial pregnancy and baby care information. Many users rely on these search results as standalone resources without seeking medical professionals validation. The potential risks to expectant parents and newborns from false or low-quality information make it essential to scrutinize the quality of search outputs on these topics.
@@
Prior work on the quality of web search results related to baby care and pregnancy is very limited and focused on few particular aspects of baby care/pregnancy such as pregnancy-related nausea, examining information quality in the “10 blue links” – websites linked on Google Search Engine Result Pages (SERPs) (Monje and Gimovsky 2023; Sacks and Abenhaim 2013; Shao, Tulandi, and Abenhaim 2020; Artieta-Pinedo et al. 2018). However, in recent years, Google Search has evolved with many new components added to SERPs  (Oliveira and Teixeira Lopes 2023), including featured snippets (FS) and AI Overviews (AIO) – visual boxes containing direct answers powered by information retrieval and generative LLMs, respectively (Figure 1).
@@
Research on information quality within these elements, however, is scarce, both for pregnancy/baby care and more generally. Despite Google’s YMYL (Your Money or Your Life) policy requiring enhanced scrutiny for health-related content (Google 2023), top-ranked results can still contain false health claims. Prior research shows this contributes to misinformation spread (Cai, King, and Dwyer 2021). Even websites rated as high-quality by work (Lin et al. 2023), like Healthline, now include AI-generated articles that may contain unreliable medical information (Dupré 2024).
Another research gap concerns how query characteristics like question type and query sentiment relate to the display and quality of AIO/FS answers. This gap is significant given prior findings on query formulation effects. Question-based queries help users achieve their goals with minimal reformulation (Vanderschantz and Hinze 2017) and trigger better direct answers on Google (Strzelecki and Rutecka 2020). Question rewriting improves QA system performance (Buck et al. 2017; Chu et al. 2020). Certain question types (e.g., where-questions) yield higher-quality answers than others (e.g., how-questions)(Zhao et al. 2019). Query sentiment presents additional concerns. Research shows that sentiment in queries can influence responses in Generative AI search engines, with responses aligning with query bias (Li and Sinnamon 2024; Venkit et al. 2024). Such bias is particularly dangerous in high-stakes domains. Consider pregnancy-related queries like “is it safe/unsafe for pregnant women to take [medicine/food].” Answers that mistake “safe” for “unsafe” may cause severe health consequences. Therefore, high-stakes health information should not depend on query sentiment polarity. These concerns, combined with the above evidence, highlight the critical need to examine how question type and sentiment influence both the appearance and quality of AI-driven components like AIO and FS in high-stakes domains such as baby care and pregnancy.
@@
Related Work
Featured Snippet Measurement/Audit
Featured snippets, which are pieces of information extracted from webpages and highlighted at the top of search results (prior to Google AIO’s arrival), aim to directly address users’ information-seeking needs (Bink, Zimmerman, and Elsweiler 2022; Sullivan 2018). They are typically paired with highlighted text from the source page. Prior work shows that the most common source domain for generating featured snippets is Wikipedia, and they are often triggered by question queries (Strzelecki and Rutecka 2019, 2020).
On one hand, this feature has been shown to attract more attention and increase users’ dwell time on the search engine result page (SERP) (Wu et al. 2020; Gleason et al. 2023), triggering “good abandonment” – defined as when users’ information needs are satisfied without clicking on a SERP result (Li, Huffman, and Tokuda 2009; Williams et al. 2016). On the other hand, their presence can lead users to overestimate the trustworthiness of the information (Bink, Zimmerman, and Elsweiler 2022), and may contribute to health anxiety in health-related searches, particularly when the displayed results are distressing (Bink, Zimmerman, and Elsweiler 2022). This is especially concerning when featured snippets may spread misleading information in domains such as health or political information seeking (Lurie and Mulligan 2021; Scull 2020).
@@
Generative AI search engine audit
@@
Studies specifically auditing and evaluating generative AI search engines are limited, with few early explorations conducting evaluation via either fully manual annotation (Li and Aral 2025; Liu, Zhang, and Liang 2023; Hu et al. 2024) or hybrid methods combining manual effort with LLM-as-a-judge or machine learning classifiers (Venkit et al. 2024; Li and Sinnamon 2024). Most focus on specific answer quality properties (e.g., verifiability, sentiment, source category) or scenarios (e.g., adversarial attacks). For example, prior work (Liu, Zhang, and Liang 2023) examined verifiability through broad metrics (fluency, perceived utility) for measuring answers’ text generally and specific metrics (citation recall, citation precision) for checking answers’ reference, finding that generative AI responses often exhibit high fluency and perceived utility but frequently contain unsupported statements or inaccurate citations. Another study (Hu et al. 2024) evaluated generative search engines under adversarial input scenarios, demonstrating effectiveness in inducing incorrect responses via manipulative questions. Research in Li and Sinnamon (2024) examined sentiment in queries and answers, as well as source types, finding evidence of response sentiment bias driven by query sentiment, plus commercial and geographic biases in source categories. Research in Li and Aral (2025) specifically studied people’s trust in GenAI search engines compared to traditional ones via a preregistered, randomized experiment on a large U.S. representative sample.
@@
Among these works, Venkit et al. (2024) conducted a more comprehensive evaluation spanning answer text, citations, sources, and user interface via hybrid methods including LLM-as-a-judge and manual evaluation (user interviews). They found that GenAI answers tend to use overtly confident language and align with biases implied in questions (“confirmation bias”). However, while confirmation bias wasn’t specifically studied during implementation in (Venkit et al. 2024), and (Li and Sinnamon 2024) did not control for specific topics or question types – factors known to affect question-answering systems (Li and Aral 2025; Zhao et al. 2019; Buck et al. 2017; Chu et al. 2020) – we investigate confirmation bias more rigorously by controlling query topic and question type (RQ4). We also integrate metrics from prior studies relevant to our research setting, which we elaborate on in the following methodology section. Previous work has raised concerns about biases in LLM-as-a-judge evaluation (Ye et al. 2024) and documented only moderate correlation between human annotation and LLM-as-a-judge (e.g., Pearson correlation of 0.62 in similar search settings (Venkit et al. 2024)), while recent evidence has documented systematic risks and unreliability when using LLMs for annotation tasks without rigorous validation (Baumann et al. 2025). Given these concerns, we opt for full manual evaluation for all metrics except source category in this study.
Methodology
@@
Search Query Selection and Reformulation
@@
For the first step, we selected the queries related to baby care and pregnancy from the 10.4 million distinct Bing search queries in the Open Resource for Click Analysis in Search (ORCAS) (which is a public click-based dataset associated with information retrieval tasks [^1] This search query dataset contains no personally identifiable information and has been filtered using a k-anonymity threshold to exclude infrequent queries. Offensive and pornographic content has been removed, meaning some extreme queries may be absent.(Craswell et al. 2020). Using this dataset allows us to include real search queries actually utilized by the users when seeking information or advice on baby care and pregnancy-related questions. For the first filtering step, we used the keywords “baby” and “babies” for the topic baby care and “pregnant” and “pregnancy” for the topic pregnancy. For the second step, we filtered for queries with a minimum length of 3 words, as 2-word queries (e.g., “baby teeth,” “baby eyes”) tend to be too vague. Additionally, we removed almost identical queries. To do this, we applied tokenization, stemming, and stopwords-removal to the queries, and then removed the queries that were duplicates after this pre-processing step. As a result, we obtained a total of 15,500 queries.
In the third step, we conducted further filtering of the queries to ensure we include only those related to human baby care and/or pregnancy. This was necessary to filter out queries related to (1) baby care/pregnancy of an animal or plant [^2] We filtered out any plant-related queries that might use pregnancy-related terminology.; (2) information about pregnancy of a celebrity, fictional character, or a specific person; (3) a name, song, or item that simply contains the words “baby” or “pregnancy” but is not related to seeking advice or information on pregnancy. For this step, we relied on annotation using gpt-4o-mini. Following best practices for LLM-based annotation (Baumann et al. 2025), we validated the model’s performance on a subset of 50 manually annotated queries before applying it to our full dataset. The gpt-4o-mini classifications overlapped with our manual annotations 100% of the time. The exact prompt we used can be found in the Appendix. In total, after this step, we retained 9516 queries relevant to questions about human baby care and/or pregnancy.
@@
We observed that Google’s AI Overviews (AIO) are sometimes present in the HTML of a results page but not shown to users, and the corresponding result pages either show a banner $\star$ An AI Overview is not available for this search” on the very top of the page or nothing. The fact that Google restricts triggering visible AIO for certain queries has previously been described by  (Williams 2024). To capture this, we design a parser to identify both the visible AIOs and those embedded but hidden from view. Specifically, we first verify AIO presence by checking for both a primary identifier element and a container element. Once confirmed, we then classify AIOs as visible or suppressed based on styling attributes and the absence of certain elements. We mark AIO as “suppressed” when it lacks elements that typically accompany visible overviews and also lacks certain styling attributes, indicating that the AIO answer exists in the underlying HTML DOM structure but is deliberately hidden from users. [^3] We will release the code and data on GitHub upon publication.
Manual Evaluation Metrics (RQ2, 3, 4)
To evaluate the quality of outputs, we relied on manual annotation of the results. Building on prior work on information quality and contradictory text identification (Es et al. 2024; Saad-Falcon et al. 2023; Li, Raheja, and Kumar 2023; De Marneffe, Rafferty, and Manning 2008; Wu, Niu, and Rahman 2022; Amugongo et al. 2024), we defined five core metrics tailored to our setting, i.e., consistency, relevance, safeguard, domain category, and sentiment valence.
@@
Instead of adopting a strict logical definition of consistency, where prior work defines properties such as transitivity, commutativity, and negation invariance (Liu et al. 2024; Ghosh et al. 2024), we propose a looser definition that more closely matches human intuitions and facilitates annotation. This is inspired by prior work on contradiction detection in text (Li, Raheja, and Kumar 2023; De Marneffe, Rafferty, and Manning 2008). Given a query, consistency is defined between a pair of answers when there is no clear countervailing and contradictory statement/evidence. Namely, consistent answers provide equivalent or similar information, where one answer might offer complementary details without creating contradictions or express the same core information in different words.
In contrast, contradiction occurs when there is clear countervailing evidence between pairwise statements given a question. This can manifest through relatively obvious features (antonymy, negation, or numeric mismatches) or through complex differences in assertion structure, world-knowledge discrepancies, and lexical contrasts, as indicated in prior work (De Marneffe, Rafferty, and Manning 2008). Prior work summarizes different types of contradiction (Li, Raheja, and Kumar 2023; Wu, Niu, and Rahman 2022). We provide a taxonomy of contradictions that are tailored to our research context as follows:
  • Binary Contradictions: Direct opposites where answers are mutually exclusive (yes/no, safe/unsafe, do/don’t, etc.) and cannot both be true or co-existing simultaneously (e.g., not dangerous vs. concerning). These are clear-cut disagreements, including the negation and antonymy types defined in previous work (Li, Raheja, and Kumar 2023; Wu, Niu, and Rahman 2022).
  • Numeric Mismatches: Different, but specific ranges, times, measurements or likelihoods that don’t overlap and can create confusion; any numerical or temporal or likelihood differences or mismatches, following a definition in prior work (Li, Raheja, and Kumar 2023);
@@
In addition to comparing the complete AIO and FS answers, we separately evaluated the consistency between their highlighted portions. This dual assessment simulates two distinct user behaviors: the “quick scanner” who only reads highlighted text before drawing conclusions, and the “thorough reader” who examines the entire content of both components. This approach was inspired by Fernández-Pichel, Pichel, and Losada (2024), who similarly modeled different user behavior modes (lazy versus diligent users) during health information seeking.
@@
Fortiguard categorizes websites but does not reliably indicate source quality: for instance, websites classified as Health can include both highly reliable scientific medical websites and sources such as Healthline that are effectively commercial media and contain unreliable, AI-generated content (Dupré 2024). Thus, in a second step, we manually classified 144 unique domains (i.e., the top 10% websites most frequently mentioned in AIO, FS, and 10 blue links as: High credibility sources, including government and international health institutions (e.g., CDC, NHS, WHO), hospitals/clinics, medical associations (e.g., ACOG), and peer-reviewed scientific publications. Medium/contextually-dependent credibility sources, comprised of health-focused commercial media, healthcare service providers, general commercial media, non-profits, and reference websites like Wikipedia. Low credibility sources, including general and health-related e-commerce sites and social media platforms. The categorization was conducted by two of the authors who worked collaboratively to resolve all disagreements through discussion until consensus was reached.
@@
The consistency category was manually annotated by 3 coders, the corresponding Krippendorff’s Alpha was 0.74 on 100 samples. . All other categories were manually annotated by 2 coders, with the Krippendorff’s Alphas as follows: 0.98 for relevance on 100 samples, 0.81 for sentiment annotation on 33 sample and 0.84 for safeguard annotation on 30 sample samples. Thus, across all categories we reached moderate (inconsistency) to high (all other categories) levels of agreement. Given that inconsistency involved 3 annotators and is, in general, a more nuanced metric, our level of agreement is acceptable for drawing further conclusions. All annotators discussed disagreements before annotating the remaining data.
Results
RQ1: Prevalence of AIO and FS

### v3

@@ Abstract
Google Search increasingly surfaces AI-generated content through features like AI Overviews (AIO) and Featured Snippets (FS), which users frequently rely on despite having no control over their presentation. Through a systematic algorithm audit of 1,508 real baby care and pregnancy-related queries, we evaluate the quality and consistency of these information displays. Our robust evaluation framework assesses multiple quality dimensions, including answer consistency, relevance, presence of medical safeguards, source categories, and sentiment alignment. Our results reveal concerning gaps in information consistency, with information in AIO and FS displayed on the same search result page being inconsistent with each other in 33% of cases. Despite high relevance scores, both features critically lack medical safeguards (present in just 11% of AIO and 7% of FS responses). While health and wellness websites dominate the source categories for both AIO and FS, FS more often link to commercial sources. These findings have important implications for public health information access and demonstrate the need for stronger quality controls in AI-mediated health information. Our methodology provides a transferable framework for auditing AI systems across high-stakes domains where information quality directly impacts user well-being.
@@
Figure 1: Overview of our audit study methods and results on Google’s AI Overviews (AIO) and Featured Snippets (FS) for pregnancy and baby care information. In the example screenshot, the AIO appears above the FS. The first sentence of the AIO is highlighted by Google to add emphasis. We find that AIO occur more frequently than FS (RQ1) and show considerable inconsistency with FS answers, more profound in highlighted pairs (RQ2). While AIO and FS responses are generally relevant (RQ3a), they provide safeguard cues infrequently (RQ3b), and FS sources concern significantly higher proportions from commercial categories than AIO or ten blue links results (RQ3c). RQ4: We do not find evidence of “confirmation bias” in AIO answers, where a user’s sentiment (positive/neutral/negative) is reflected in the query.
1 Introduction
@@
Search engines like Google serve as critical information gatekeepers for high-stakes topics where people need to rapidly acquire extensive knowledge to make important decisions. This includes domains like healthcare, finance, and politics, where search results significantly impact individual and societal well-being (Fox 2011; Fernández-Pichel et al. 2024). Among these areas, baby care and pregnancy offer a compelling example – new and expectant parents make frequent, time-sensitive decisions while learning from many new concepts, often relying on search engines during a critical period lasting months or years. Specifically, 85% of pregnant women surveyed by (Giacometti et al. 2024) turned to Google for pregnancy-related information in 2024. While over 80% stated that the information they found online had a medium to very high influence on their decision-making related to pregnancy, only 42% reported validating such information with their doctors (Giacometti et al. 2024). Thus, web search, particularly Google, has become a primary source for crucial pregnancy and baby care information. Many users rely on these search results as standalone resources without seeking medical professionals’ validation. The potential risks to expectant parents and newborns from false or low-quality information make it essential to scrutinize the quality of search outputs on these topics.
@@
Prior work on the quality of web search results related to baby care and pregnancy is very limited and focused on few particular aspects of baby care/pregnancy such as pregnancy-related nausea, examining information quality in the “ten blue links” – websites linked on Google Search Engine Result Pages (SERPs) (Monje and Gimovsky 2023; Sacks and Abenhaim 2013; Artieta-Pinedo et al. 2018). However, in recent years, Google Search has evolved with many new components added to SERPs (Oliveira and Teixeira Lopes 2023), including featured snippets (FS) and AI Overviews (AIO) – visual boxes containing direct answers powered by information retrieval and generative LLMs, respectively (Figure 1).
@@
Research on information quality within these elements, however, is scarce, both for pregnancy/baby care and more generally. Despite Google’s YMYL (Your Money or Your Life) policy requiring enhanced scrutiny for health-related content (Google 2023), top-ranked results can still contain false health claims. Prior research shows this contributes to misinformation spread (Cai et al. 2021). Even websites rated as high-quality by work (Lin et al. 2023), like Healthline, now include AI-generated articles that may contain unreliable medical information (Dupré 2024).
Another research gap concerns how query characteristics like question type and query sentiment relate to the display and quality of AIO/FS answers. This gap is significant given prior findings on query formulation effects. Question-based queries help users achieve their goals with minimal reformulation (Vanderschantz and Hinze 2017) and trigger better direct answers on Google (Strzelecki and Rutecka 2020). Question rewriting improves QA system performance (Buck et al. 2017; Chu et al. 2020). Certain question types (e.g., where-questions) yield higher-quality answers than others (e.g., how-questions)(Zhao et al. 2019). Query sentiment presents additional concerns. Research shows that sentiment in queries can influence responses in Generative AI search engines, with responses aligning with query bias (Venkit et al. 2024). Such bias is particularly dangerous in high-stakes domains (Li and Sinnamon 2024; Baumann et al. 2025b). Consider pregnancy-related queries like “is it safe/unsafe for pregnant women to take [medicine/food].” Answers that mistake “safe” for “unsafe” may cause severe health consequences. Therefore, high-stakes health information should not depend on query sentiment polarity. These concerns, combined with the above evidence, highlight the critical need to examine how question type and sentiment influence both the appearance and quality of AI-driven components like AIO and FS in high-stakes domains such as baby care and pregnancy.
@@
2 Related Work
2.1 Featured Snippet Measurement/Audit
Featured snippets, which are pieces of information extracted from webpages and highlighted at the top of search results (prior to Google’s introduction of AIO), aim to directly address users’ information-seeking needs (Bink et al. 2022). They are typically paired with highlighted text from the source page. Prior work shows that the most common source domain for generating featured snippets is Wikipedia, and they are often triggered by question queries (Strzelecki and Rutecka 2019, 2020).
On one hand, this feature has been shown to attract more attention and increase users’ dwell time on the search engine result page (SERP) (Wu et al. 2020; Gleason et al. 2023), triggering “good abandonment” – defined as when users’ information needs are satisfied without clicking on a SERP result (Li et al. 2009). On the other hand, their presence can lead users to overestimate the trustworthiness of the information (Bink et al. 2022), and may contribute to health anxiety in health-related searches, particularly when the displayed results are distressing (Bink et al. 2022). This is especially concerning when featured snippets may spread misleading information in domains such as health or political information seeking (Lurie and Mulligan 2021; Scull 2020).
@@ 2.1 Featured Snippet Measurement/Audit
2.2 Generative AI search engine audit
@@
Studies specifically auditing and evaluating generative AI search engines are limited, with few early explorations conducting evaluation via either fully manual annotation (Li and Aral 2025; Liu et al. 2023; Hu et al. 2024) or hybrid methods combining manual effort with LLM-as-a-judge or machine learning classifiers (Venkit et al. 2024; Li and Sinnamon 2024). Most focus on specific answer quality properties (e.g., verifiability, sentiment, source category) or scenarios (e.g., adversarial attacks). For example, prior work (Liu et al. 2023) examined verifiability through broad metrics (fluency, perceived utility) for measuring answers’ text generally and specific metrics (citation recall, citation precision) for checking answers’ reference, finding that generative AI responses often exhibit high fluency and perceived utility but frequently contain unsupported statements or inaccurate citations. Another study (Hu et al. 2024) evaluated generative search engines under adversarial input scenarios, demonstrating effectiveness in inducing incorrect responses via manipulative questions. Research in Li and Sinnamon (2024) examined sentiment in queries and answers, as well as source types, finding evidence of response sentiment bias driven by query sentiment, plus commercial and geographic biases in source categories. Research in Li and Aral (2025) specifically studied people’s trust in GenAI search engines compared to traditional ones via a preregistered, randomized experiment on a large U.S. representative sample.
@@
Among these works, Venkit et al. (2024) conducted a more comprehensive evaluation spanning answer text, citations, sources, and user interface via hybrid methods including LLM-as-a-judge and manual evaluation (user interviews). They found that GenAI answers tend to use overtly confident language and align with biases implied in questions (“confirmation bias”). However, while confirmation bias wasn’t specifically studied during implementation in (Venkit et al. 2024), and (Li and Sinnamon 2024) did not control for specific topics or question types – factors known to affect question-answering systems (Li and Aral 2025; Zhao et al. 2019; Buck et al. 2017; Chu et al. 2020) – we investigate confirmation bias more rigorously by controlling query topic and question type (RQ4). We also integrate metrics from prior studies relevant to our research setting, which we elaborate on in the following methodology section. Previous work has raised concerns about biases in LLM-as-a-judge evaluation (Ye et al. 2024) and documented only moderate correlation between human annotation and LLM-as-a-judge (e.g., Pearson correlation of 0.62 in similar search settings (Venkit et al. 2024)), while recent evidence has documented systematic risks and unreliability when using LLMs for annotation tasks without rigorous validation (Baumann et al. 2025a). Given these concerns, we opt for full manual evaluation for all metrics except source category in this study.
3 Methodology
@@
3.1 Search Query Selection and Reformulation
@@
For the first step, we selected the queries related to baby care and pregnancy from the 10.4 million distinct Bing search queries in the Open Resource for Click Analysis in Search (ORCAS), which is a public click-based dataset associated with information retrieval tasks [^1] This search query dataset contains no personally identifiable information and has been filtered using a k-anonymity threshold to exclude infrequent queries. Offensive and pornographic content has been removed, meaning some extreme queries may be absent. (Craswell et al. 2020). Queries on baby care and pregnancy represent persistent everyday information needs rather than viral spikes, appearing in only 0.05% of trend episodes in the GoogleTrendArchive dataset (Urman et al. 2026). In contrast, ORCAS captures the long tail of real user queries, making it better suited for studying realistic health information-seeking behavior.
For the first filtering step, we used the keywords “baby” and “babies” for the topic baby care and “pregnant” and “pregnancy” for the topic pregnancy. For the second step, we filtered for queries with a minimum length of 3 words, as 2-word queries (e.g., “baby teeth,” “baby eyes”) tend to be too vague. Additionally, we removed almost identical queries. To do this, we applied tokenization, stemming, and stopwords-removal to the queries, and then removed the queries that were duplicates after this pre-processing step. As a result, we obtained a total of 15,500 queries.
In the third step, we conducted further filtering of the queries to ensure we include only those related to human baby care and/or pregnancy. This was necessary to filter out queries related to (1) baby care/pregnancy of an animal or plant [^2] We filtered out any plant-related queries that might use pregnancy-related terminology.; (2) information about pregnancy of a celebrity, fictional character, or a specific person; (3) a name, song, or item that simply contains the words “baby” or “pregnancy” but is not related to seeking advice or information on pregnancy. For this step, we relied on annotation using gpt-4o-mini. Following best practices for LLM-based annotation (Baumann et al. 2025a), we validated the model’s performance on a subset of 50 manually annotated queries before applying it to our full dataset. The gpt-4o-mini classifications overlapped with our manual annotations 100% of the time. The exact prompt we used can be found in the Appendix. In total, after this step, we retained 9516 queries relevant to questions about human baby care and/or pregnancy.
@@ 3.2 Auditing Pipeline
We observed that Google’s AI Overviews (AIO) are sometimes present in the HTML of a results page but not shown to users, and the corresponding result pages either show a banner $\star$ An AI Overview is not available for this search” on the very top of the page or nothing. The fact that Google restricts triggering visible AIO for certain queries has previously been described by (Williams 2024). To capture this, we design a parser to identify both the visible AIOs and those embedded but hidden from view. Specifically, we first verify AIO presence by checking for both a primary identifier element and a container element. Once confirmed, we then classify AIOs as visible or suppressed based on styling attributes and the absence of certain elements. We mark AIO as “suppressed” when it lacks elements that typically accompany visible overviews and also lacks certain styling attributes, indicating that the AIO answer exists in the underlying HTML DOM structure but is deliberately hidden from users. [^3] Code and data are available at https://github.com/whocheers/Google-AIO-FS-Audit.
3.3 Manual Evaluation Metrics (RQ2, 3, 4)
To evaluate the quality of outputs, we relied on manual annotation of the results. Building on prior work on information quality and contradictory text identification (Es et al. 2024; Saad-Falcon et al. 2023; Li et al. 2023; De Marneffe et al. 2008; Wu et al. 2022), we defined five core metrics tailored to our setting, i.e., consistency, relevance, safeguard, domain category, and sentiment valence.
@@ 3.3 Manual Evaluation Metrics (RQ2, 3, 4)
Instead of adopting a strict logical definition of consistency, where prior work defines properties such as transitivity, commutativity, and negation invariance (Liu et al. 2024), we propose a looser definition that more closely matches human intuitions and facilitates annotation. This is inspired by prior work on contradiction detection in text (Li et al. 2023; De Marneffe et al. 2008). Given a query, consistency is defined between a pair of answers when there is no clear countervailing and contradictory statement/evidence. Namely, consistent answers provide equivalent or similar information, where one answer might offer complementary details without creating contradictions or express the same core information in different words.
In contrast, contradiction occurs when there is clear countervailing evidence between pairwise statements given a question. This can manifest through relatively obvious features (antonymy, negation, or numeric mismatches) or through complex differences in assertion structure, world-knowledge discrepancies, and lexical contrasts, as indicated in prior work (De Marneffe et al. 2008). Prior work summarizes different types of contradiction (Li et al. 2023; Wu et al. 2022). We provide a taxonomy of contradictions that are tailored to our research context as follows:
  • Binary Contradictions: Direct opposites where answers are mutually exclusive (yes/no, safe/unsafe, do/don’t, etc.) and cannot both be true or co-existing simultaneously (e.g., not dangerous vs. concerning). These are clear-cut disagreements, including the negation and antonymy types defined in previous work (Li et al. 2023; Wu et al. 2022).
  • Numeric Mismatches: Different, but specific ranges, times, measurements or likelihoods that don’t overlap and can create confusion; any numerical or temporal or likelihood differences or mismatches, following a definition in prior work (Li et al. 2023);
@@
In addition to comparing the complete AIO and FS answers, we separately evaluated the consistency between their highlighted portions. This dual assessment simulates two distinct user behaviors: the “quick scanner” who only reads highlighted text before drawing conclusions, and the “thorough reader” who examines the entire content of both components. This approach was inspired by Fernández-Pichel et al. (2024), who similarly modeled different user behavior modes (lazy versus diligent users) during health information seeking.
@@
Fortiguard categorizes websites but does not reliably indicate source quality: for instance, websites classified as Health can include both highly reliable scientific medical websites and sources such as Healthline that are effectively commercial media and contain unreliable, AI-generated content (Dupré 2024). Thus, in a second step, we manually classified 144 unique domains (i.e., the top 10% websites most frequently mentioned in AIO, FS, and ten blue links as: High credibility sources, including government and international health institutions (e.g., CDC, NHS, WHO), hospitals/clinics, medical associations (e.g., ACOG), and peer-reviewed scientific publications. Medium/contextually-dependent credibility sources, comprised of health-focused commercial media, healthcare service providers, general commercial media, non-profits, and reference websites like Wikipedia. Low credibility sources, including general and health-related e-commerce sites and social media platforms. The categorization was conducted by two of the authors who worked collaboratively to resolve all disagreements through discussion until consensus was reached.
@@
The consistency category was manually annotated by 3 coders, the corresponding Krippendorff’s Alpha was 0.74 on 100 samples. All other categories were manually annotated by 2 coders, with the Krippendorff’s Alphas as follows: 0.98 for relevance on 100 samples, 0.81 for sentiment annotation on 33 sample and 0.84 for safeguard annotation on 30 sample samples. Thus, across all categories we reached moderate (inconsistency) to high (all other categories) levels of agreement. Given that inconsistency involved 3 annotators and is, in general, a more nuanced metric, our level of agreement is acceptable for drawing further conclusions. All annotators discussed disagreements before annotating the remaining data.
4 Results
4.1 RQ1: Prevalence of AIO and FS